OpenStudy (anonymous):

This one is probably a bit interesting: Suppose \( K\) be the number of integers \( n \) such that \( \large \frac{2^n+1}{n^2}\) is also an integer.Find \(K\). PS:This was posted (by me) earlier in Mathematics group, I am cross-posting it here as this is more meta material.

OpenStudy (binary3i):

not sure but is K=2

OpenStudy (anonymous):

Yes, but how about proving it ? ;)

OpenStudy (binary3i):

the graph(2^n+1=y) and the graph (An^2=y) intersect at only two points.

OpenStudy (anonymous):

So you have plotted through out infinity ? and graphical proof isn't admissible in number theory.

OpenStudy (binary3i):

ok then let me try

OpenStudy (binary3i):

2^n=(An^2-1) take log. nlog2=log(An^2-1) take its derivative. log2=1/(An^2-1) (A2n) gives a qudratic which has two values of n.

OpenStudy (binary3i):

what do you think is it correct?

OpenStudy (anonymous):

can you please explain question properly i didnt get what is relation between n ,k

OpenStudy (anonymous):

IMO-1990-problem 3